# Neural Networks for Extreme Quantile Regression with an Application to Forecasting of Flood Risk

### v3

@@ 6 Conclusion
Our EQRN model combines extrapolation results from extreme value theory with the prediction power of neural networks. It provides a flexible and versatile method for extreme quantile regression that is capable of prediction beyond the range of the data in the presence of a large number of covariates. The main focus in this paper was the case of sequential dependence to develop a tool for risk forecasting in time series that can be used for effective early warning systems in flood management. Our model already performs well in issuing sparse warnings for the days with increased risk of flooding, as illustrated in the case study of the Aare catchment in Switzerland. A further improvement could be attained by using additional covariates as input of the model. This could include observations of variables that are typically used in hydrological models such as soil moisture, or forecasts of atmospheric variables such as precipitation and temperature.
Many other applications seem pertinent. Even in the case of independent data, our simulation study in Supplementary Material S.3 shows that neural networks outperform tree-based methods such as ERF and GBEX if the quantile function is more complex. Applications are financial risk assessment in insurance companies or banks. For spatial data, images or graphs, convolutional or graph neural networks (LeCun et al., 2015; Scarselli et al., 2009; Wu et al., 2021) are known to perform extremely well in capturing neighbourhood structures. Our EQRN method can therefore be applied to quantify the risk of climate extremes where the predictor space contains spatio-temporal observations of meteorological variables (e.g., Boulaguiem et al., 2022). Transformer architectures (Vaswani et al., 2017) can also perform well for spatio-temporal or more complex dependencies.
The price for the high flexibility of machine learning methods, which focus on prediction accuracy, is limited statistical interpretability. However, feature-importance identification methods are becoming increasingly popular for interpreting neural network predictions (Lundberg and Lee, 2017). There is also active research on the construction of prediction intervals for black-box methods, for instance through conformal inference (Lei et al., 2018; Romano et al., 2019). How such techniques can be adapted to assess uncertainty for extreme quantile regression is an interesting future research question.
@@ Acknowledgements
The authors would like to thank Daniel Viviroli for his valuable insights.
@@
Figure S.1 shows a schematic representation of a multi-layer LSTM network.
@@
  1. $\mathcal{I}$): For independent data, splits the index set $\mathcal{I}$ randomly into training set $\mathcal{T}$ and validation set $\mathcal{V}$ with pre-specified proportions.
  2. $\mathcal{I}$): For sequential data, splits the index set $\mathcal{I}$ sequentially into training set $\mathcal{T}$ and validation set $\mathcal{V}$ with pre-specified proportions, such that all observations in $\mathcal{T}$ are before $\mathcal{V}$ in time.
@@
Figure S.1: Multi-layer LSTM network flowchart from input $\tilde{\mathbf{x}}_{t}:=(\mathbf{x}_{t-s},\ldots,\mathbf{x}_{t-1})$ to output $\tilde{g}_{\mathcal{W}}(\tilde{\mathbf{x}}_{t})$, with loss evaluation. The LSTM cells represent the transformation in (9).
@@
Figure S.2 shows the accuracy of EQRN and the competitor methods for an increasingly large $\tau$, for the three data models. The competitors’ performances shown here were also significantly improved by using the intermediate $\hat{Q}_{{\mathbf{x}}}(\tau_{0})$ as an additional covariate. For every model, EQRN outperforms all competitors, with a difference in accuracy increasing with $\tau$. EGAM seems to suffer from the large dimension of the feature space at large $\tau$, both when the actual quantile depends on only two or all of the ten features. The difference in accuracy of EQRN and GBEX compared to the unconditional and semi-conditional models is particularly significant for Models 1 and 3, and EQRN generally outperforms GBEX, especially at high probability levels.
@@
Figure S.4 shows the predicted $\hat{Q}_{\mathbf{x}}(0.9995)$ for EQRN and the competitors, as a function of the two significant covariates $(x_{1},x_{2})$ for Model 1. At that extreme level, EQRN still seems to capture the true conditional quantile function quite well, although the predictions show some residual noise. GBEX shows an elliptical stepwise approximation behaviour. and seems to underestimate the smaller quantiles and overestimate the largest quantiles. EGAM and GRF fail drastically at recovering the conditional quantile function. The semi-conditional estimates are a translation of the intermediate quantiles, which in particular fail to capture the varying shape along $X_{1}$.
@@
Figure S.7 shows the impact of the intermediate level $\tau_{0}$ on the accuracy of the best EQRN model. The value $\tau_{0}=0.8$ used in the rest of the analysis leads to the best RMSE. This relatively low value shows that $\tau_{0}$ can be chosen much lower than for the classical unconditional GPD model. An intuitive explanation for this fact is the following. In a situation without covariates, the choice of $\tau_{0}$ is a trade-off between approximation bias (which favours larger thresholds) and variance (which favours lower thresholds); see Figure 3 in the main document. For covariate-dependent data, the distribution of the exceedances varies and more data is needed to accurately capture this function of the covariates. The variance becomes more important than the approximation bias, and therefore lower thresholds are preferable. Moreover, the flexibility of the GPD regression neural network model seems to be able to absorb some of the approximation bias, also allowing for a low value of $\tau_{0}$.
@@
We first observe that the predictions have roughly the same behaviour, which shows that all methods capture at least some of the temporal structure based on the past covariates. The semi-conditional method (Figure S.8) is clearly not flexible enough, since the predictions only show a weak sensitivity to the changes in covariates. It also fails to trigger any early warning during the main event, due to low probability ratio forecasts never exceeding the selected threshold value of $100$. The reason is that, while the intermediate quantile is covariate-dependent, the GPD parameters are constant over time. We conclude that a covariate-dependent model for the tail is required in this application. Figure S.9 shows predictions from the EXQAR model (Li and Wang, 2019). This model is more sensitive to changes in the covariates, but the regression function looks fairly erratic, with sudden spikes at some time points. Those spikes are here due to unusually small shape estimates in combination with a large-scale estimate. This might be caused by an instability in the estimation of the local moments used in the model. EGAM (Figure S.10) fails to trigger an early warning for the first day of the flooding event as its quantile and probability ratio forecasts are very low, but then seems to severely overestimate the river flow during the rest of the event. The best competing model seems to be the GBEX (Figure S.11), as it yields a smooth prediction curve with a pronounced spike at the main event. Comparing this with our EQRN method (Figure 2 in the main paper), it reacts slightly later and its risk forecast (probability ratio) on the day before the first exceedance is significantly lower than for EQRN.
Figure S.8: Top: Daily average discharge (points) at the Bern–Schönau station (62) and one-day-ahead semi-conditional forecasts of conditional 100-year quantiles (solid line) during the 2005 flood. Horizontal lines show unconditional $Q^{100}$ based on GEV (dashed) and GPD (dotted). Bottom: one-day-ahead forecast of the conditional probability of exceeding the GEV estimated $Q^{100}$ as a ratio to the unconditional probability, using the semi-conditional parameter forecast. The vertical line indicates August 22, the day of the first exceedance.

### v4

@@ 6 Conclusion
Our EQRN model combines extrapolation results from extreme value theory with the prediction power of neural networks. It provides a flexible and versatile method for extreme quantile regression that is capable of prediction beyond the range of the data in the presence of a large number of covariates. Customised network architectures can be used in our open-source “$\mathtt{EQRN}$” R package, allowing for tailor-made models capturing all types of potential dependencies between covariates and between observations.
The main focus in this paper was the case of sequential dependence to develop a tool for risk forecasting in time series that can be used for effective early warning systems in flood management. Our model already performs well in issuing sparse warnings for the days with increased risk of flooding, as illustrated in the case study of the Aare catchment in Switzerland. A further improvement could be attained by using additional covariates as input of the model. This could include observations of variables that are typically used in hydrological models, such as soil moisture, or forecasts of atmospheric variables such as precipitation and temperature.
Many other applications seem pertinent. Even in the case of independent data, our simulation study in Supplementary Material S.3 shows that neural networks outperform tree-based methods, such as ERF and GBEX, if the quantile function is more complex. Applications are financial risk assessment in insurance companies or banks. For spatial data, images, or graphs, convolutional or graph neural networks (LeCun et al., 2015; Scarselli et al., 2009; Wu et al., 2021) are known to perform extremely well in capturing neighbourhood structures. Our EQRN method can, therefore, be applied to quantify the risk of climate extremes where the predictor space contains spatiotemporal observations of meteorological variables (e.g., Boulaguiem et al., 2022). Transformer architectures (Vaswani et al., 2017) can also perform well for spatiotemporal or more complex dependencies.
The price for the high flexibility of machine learning methods, which focus on prediction accuracy, is limited statistical interpretability. However, feature-importance identification methods are becoming increasingly popular for interpreting neural network predictions (Lundberg and Lee, 2017). There is also active research on the construction of prediction intervals for black-box methods, for instance, through conformal inference (Lei et al., 2018; Romano et al., 2019). How such techniques can be adapted to assess uncertainty for extreme quantile regression is an interesting future research question.
Declarations
@@ Acknowledgements
The authors would like to thank Daniel Viviroli for his valuable insights as well as the editorial team of the Annals of Applied Statistics and the anonymous reviewers for their helpful comments.
@@
Published article
This document is the peer-reviewed “Author’s Accepted Manuscript” of an article published in the Annals of Applied Statistics (Pasche and Engelke, 2024), with the DOI https://doi.org/10.1214/24-AOAS1907. When citing this work, please refer to the published version.
@@
Figure S.1 shows a schematic representation of a multilayer LSTM network.
@@
  1. $\mathcal{I}$): For independent data, splits the index set $\mathcal{I}$ randomly into training set $\mathcal{T}$ and validation set $\mathcal{V}$ with prespecified proportions.
  2. $\mathcal{I}$): For sequential data, splits the index set $\mathcal{I}$ sequentially into training set $\mathcal{T}$ and validation set $\mathcal{V}$ with prespecified proportions, such that all observations in $\mathcal{T}$ are before $\mathcal{V}$ in time.
@@
Figure S.1: Multilayer LSTM network flowchart from input $\tilde{\mathbf{x}}_{t}:=(\mathbf{x}_{t-s},\ldots,\mathbf{x}_{t-1})$ to output $\tilde{g}_{\mathcal{W}}(\tilde{\mathbf{x}}_{t})$, with loss evaluation. The LSTM cells represent the transformation in (9).
@@
Figure S.2 shows the accuracy of EQRN and the competitor methods for an increasingly large $\tau$, for the three data models. The competitors’ performances shown here were also significantly improved by using the intermediate $\hat{Q}_{{\mathbf{x}}}(\tau_{0})$ as an additional covariate. For every model, EQRN outperforms all competitors, with a difference in accuracy increasing with $\tau$. EGAM seems to suffer from the large dimension of the feature space at large $\tau$, both when the actual quantile depends on only two or all of the 10 features. The difference in accuracy of EQRN and GBEX compared to the unconditional and semiconditional models is particularly significant for Models 1 and 3, and EQRN generally outperforms GBEX, especially at high probability levels.
@@
Figure S.4 shows the predicted $\hat{Q}_{\mathbf{x}}(0.9995)$ for EQRN and the competitors, as a function of the two significant covariates $(x_{1},x_{2})$ for Model 1. At that extreme level, EQRN still seems to capture the true conditional quantile function quite well, although the predictions show some residual noise. GBEX shows an elliptical stepwise approximation behaviour. and seems to underestimate the smaller quantiles and overestimate the largest quantiles. EGAM and GRF fail drastically at recovering the conditional quantile function. The semiconditional estimates are a translation of the intermediate quantiles, which in particular fail to capture the varying shape along $X_{1}$.
@@
Figure S.7 shows the impact of the intermediate level $\tau_{0}$ on the accuracy of the best EQRN model. The value $\tau_{0}=0.8$ used in the rest of the analysis leads to the best RMSE. This relatively low value shows that $\tau_{0}$ can be chosen much lower than for the classical unconditional GPD model. An intuitive explanation for this fact is the following. In a situation without covariates, the choice of $\tau_{0}$ is a trade-off between approximation bias (which favours larger thresholds) and variance (which favours lower thresholds); see Figure 3 in the main document. For covariate-dependent data, the distribution of the exceedances varies and more data is needed to accurately capture this function of the covariates. The variance becomes more important than the approximation bias, therefore, lower thresholds are preferable. Moreover, the flexibility of the GPD regression neural network model seems to be able to absorb some of the approximation bias, also allowing for a low value of $\tau_{0}$.
@@
We first observe that the predictions have roughly the same behaviour, which shows that all methods capture at least some of the temporal structure based on the past covariates. The semiconditional method (Figure S.8) is clearly not flexible enough, since the predictions only show a weak sensitivity to the changes in covariates. It also fails to trigger any early warning during the main event, due to low probability ratio forecasts never exceeding the selected threshold value of $100$. The reason is that, while the intermediate quantile is covariate-dependent, the GPD parameters are constant over time. We conclude that a covariate-dependent model for the tail is required in this application. Figure S.9 shows predictions from the EXQAR model (Li and Wang, 2019). This model is more sensitive to changes in the covariates, but the regression function looks fairly erratic, with sudden spikes at some time points. Those spikes are here due to unusually small shape estimates in combination with a large-scale estimate. This might be caused by an instability in the estimation of the local moments used in the model. EGAM (Figure S.10) fails to trigger an early warning for the first day of the flooding event as its quantile and probability ratio forecasts are very low, but then seems to severely overestimate the river flow during the rest of the event. The best competing model seems to be the GBEX (Figure S.11), as it yields a smooth prediction curve with a pronounced spike at the main event. Comparing this with our EQRN method (Figure 2 in the main paper), it reacts slightly later and its risk forecast (probability ratio) on the day before the first exceedance is significantly lower than for EQRN.
Figure S.8: Top: Daily average discharge (points) at the Bern–Schönau station (62) and one-day-ahead semiconditional forecasts of conditional 100-year quantiles (solid line) during the 2005 flood. Horizontal lines show unconditional $Q^{100}$ based on GEV (dashed) and GPD (dotted). Bottom: One-day-ahead forecast of the conditional probability of exceeding the GEV estimated $Q^{100}$ as a ratio to the unconditional probability, using the semiconditional parameter forecast. The vertical line indicates August 22, the day of the first exceedance.